The roots of aquatic equation are 4 and -5 which quadratic equation has these roots
a.(x-4)(x+5)=0
b.(x+4)(X-5)=0
c.(x-4)(x-5)=0
d.(x+4)(x+5)=0

Answer by rothauserc(4718)

x = 4 implies that x - 4 = 0
x = -5 implies that x + 5 = 0
:
therefore, the factored quadratic equation is (x-4) * (x+5) = 0
:
answer is a.
: